# Harmonia: A Multi-Agent Reinforcement Learning Approach to Data Placement and Migration in Hybrid Storage Systems

### v1

@@ 3.1. Methodology
Data-Placement Techniques. We evaluate two state-of-the-art data-placement techniques for HSS: CDE (Cold Data Eviction) (matsui2017design, ) and Sibyl (singh2022sibyl, ). CDE is a heuristic-based data-placement technique that places hot or random (cold or sequential) data in the fast (slow) storage device, by determining the hotness and randomness of incoming data based on past access frequency and request size, respectively. Sibyl (singh2022sibyl, ) uses an RL agent to place data in the best-fit storage device in the HSS based on multiple features of I/O requests (e.g., request type and size) and HSS conditions (e.g., device capacity and I/O latency). When the fast device becomes full, both techniques evict the least-recently-used (LRU) data before placing a new I/O request data to the fast device.
@@
Data-Migration Techniques. We evaluate two state-of-the-art data-migration techniques for HSS, K-SVM (shetti2019machine, ) and RNN-HSS (doudali2019kleio, ), which have three aspects in common. First, they both leverage supervised learning to identify the target storage device for migration of each stored page within the HSS based on the page’s access frequency. Second, they periodically determine migration candidates (i.e., pages) and perform data migration at fixed intervals (e.g., migration is performed after every 1000 incoming requests) in the critical path of I/O request handling. Third, they perform data placement for incoming I/O requests using a fixed heuristic policy (e.g., K-SVM places incoming requests in the fast device until its capacity reaches 100% (shetti2019machine, )). The key difference between the two techniques is in the supervised-learning method used; K-SVM uses a k-means-assisted support vector machine (SVM) to classify stored pages as belonging to a storage device based on their past access frequencies; RNN-HSS uses recurrent neural networks (RNN) to predict a page’s future access frequency, and migrate the page to its target storage device.
@@
Extended Data-Management Techniques. To our knowledge, there is no existing technique that optimizes both data placement and data migration together, so we evaluate four extended techniques to explore the design space for a holistic data-management technique. For the first two techniques, Sibyl+K-SVM and Sibyl+RNN-HSS, we simply combine Sibyl, the current best-performing data-placement technique, with the prior data-migration techniques explained above.
@@
For the next two techniques, we explore the performance of an RL-based data-migration technique when combined with data placement. We design the third extended data-management technique, CDE+RL-Migr, [^2] The RL-based data-migration technique in CDE+RL-Migr is a contribution of this work. We design this technique as there are no prior RL-based data-migration techniques for HSS. by combining CDE, a heuristic data-placement technique, with an RL-based data-migration approach. The RL-based data-migration approach uses the same design as the data-migration agent proposed in Harmonia, including the state features and the reward structure (See §4.2).
@@
In the fourth extended technique, SAPM (Single Agent for Placement and Migration), [^3] SAPM is a contribution of this work. We design SAPM to analyze the effectiveness of a single RL agent that performs both data placement and data migration. we design a single RL agent to perform both data placement and data migration. SAPM performs a placement action for every incoming I/O request, and focuses on data migration only during system idle times to avoid interference with the placement task (which is on the critical path of I/O request handling). We provide the same state features used in Harmonia (See §4.2) to SAPM and tune the hyperparameters to achieve the best performance possible with a single RL agent. We provide SAPM’s RL agent (1) an immediate reward for every placement action (based on the reward of Harmonia’s data-placement agent in Equation 2), and (2) a deferred reward for the migration actions (based on the reward structure of Harmonia’s data-migration agent in Equation 3). We provide a detailed analysis in §3.3 on why SAPM cannot effectively optimize two different tasks using a single RL agent.
@@ 3.2. Effectiveness of Prior Techniques
In this section, we conduct an experimental study to understand the effectiveness and limitations of existing data-placement and data-migration techniques for HSS. Fig. 4 shows the performance of four prior and four extended HSS data-management techniques on (a) performance-optimized and (b) cost-optimized HSS. We compare the average request latencies of the prior techniques against two ideal approaches, Fast-Only and Oracle. First, in Fast-Only, we assume that the fast storage device is large enough to accommodate the entire workload dataset. Second, Oracle (meswani2015heterogeneous, ) makes optimal data-placement and data-migration decisions based on complete knowledge of future I/O access patterns of the entire workload. We assume that Oracle performs data migration during system idle times, which does not add any latency overheads. All values in Fig. 4 are normalized to Fast-Only.
@@
We make five key observations from Fig. 4. First, all prior techniques exhibit significant performance gaps from Oracle for every tested HSS configuration and workload. The best-performing technique, Sibyl, merely achieves 49.9% (41.4%) of Oracle’s average performance in performance-optimized (cost-optimized) HSS, which strongly implies the need for holistic data management in HSS. Sibyl’s RL-based data-placement technique can adapt to workload and system changes, but its heuristic data-eviction policy on the critical path lowers its performance.
@@
Second, Sibyl outperforms all other prior techniques, improving the average performance over CDE, K-SVM, and RNN-HSS by 20.7% (14.5%), 24.7% (31.6%), and 19.3% (19.2%), respectively, in the performance-optimized (cost-optimized) HSS. CDE, K-SVM and RNN-HSS show lower performance than Sibyl because of three key reasons: (1) they use heuristic- or supervised-learning-based approaches, which cannot easily adapt to changes in workload access patterns and HSS conditions, (2) CDE performs data eviction on the critical path of I/O request handling, which contributes to 59.2% (95.7%) of the overall latency in performance-optimized (cost-optimized) HSS, and (3) RNN-HSS and K-SVM perform migrations at fixed intervals, which adds a significant overhead to I/O request tail latencies (See §6).
@@
Third, two extended data-management techniques, Sibyl+K-SVM and Sibyl+RNN-HSS, underperform compared to Sibyl. A naive combination of prior approaches performs poorly as the placement and migration policies in such a combination are oblivious to each other’s decisions. This leads to (1) frequent page movement across storage devices by the two policies, and (2) delayed data-placement policy convergence.
@@
Fourth, CDE+RL-Migr shows lower performance than Sibyl by 4.6% (15.6%) on average in performance-optimized (cost-optimized) HSS. CDE+RL-Migr shows comparable performance to Sibyl in read-intensive workloads (i.e., RocksDB, MLPerf) because the RL-based migration agent proactively migrates frequently-read data to the fast device, which Sibyl does not do. However, CDE+RL-Migr shows lower performance in write-intensive workloads (i.e., SYSTOR17) because CDE’s heuristic data-placement policy places most incoming data in the fast storage device without considering device characteristics, resulting in frequent migrations. CDE+RL-Migr’s performance suggests that an RL-based data-migration policy can benefit performance if combined with an adaptive data-placement policy.
@@
Fifth, SAPM shows lower performance compared to Sibyl by 18.8% (27.7%) on average in the performance-optimized (cost-optimized) HSS), even though SAPM uses RL for both data placement and data migration. SAPM’s limited performance is due to two key reasons. (1) SAPM’s single RL agent is unable to learn an optimal policy for data placement and data migration concurrently, and (2) SAPM cannot identify migration candidates in a timely manner because it focuses on the data-migration task only when there are no incoming I/O requests. We provide a detailed analysis on why a single RL agent cannot optimize two different tasks in §3.3.
3.3. Need for Multi-Agent RL for
HSS Data Management
We discuss how data placement and data migration are two different tasks, and describe the limitations of a single-agent RL approach in optimizing different tasks.
3.3.1. Differences between Data Placement and Data Migration.
Data placement and data migration are two different tasks in HSS data management because of three key reasons. First, while data placement writes data from the current I/O request to the best-fit storage device, data migration monitors all previously placed data to identify candidate pages for migration and their target storage devices. Second, the objective of data placement is to minimize the I/O latency for each incoming request, but data migration aims to improve the long-term HSS performance. Third, data placement is in the critical path of I/O handling and it is performed synchronously with application writes or updates. Data migration is a background process and it is performed asynchronously to the application’s I/O requests. To describe the need for data migration in addition to data placement, we provide two examples: (1) If a page currently placed in the fast device is rarely accessed, migrating it to the slow device can free up space for more frequently-accessed pages and improve the overall HSS performance. (2) If a page on the slow device is accessed frequently (due to changes in access patterns), it should be migrated to the fast device to reduce its access latency.
3.3.2. Limitations of a Single RL Agent in Learning Diverse Tasks.
We describe two limitations of a single RL agent in optimizing multiple diverse tasks. First, a single RL agent cannot effectively learn optimal policies for multiple diverse tasks. A single-agent RL approach to optimize multiple tasks relies on task similarity (ammar2014automated, ; carroll2005task, ) to transfer the learning gained from one task to another. Task similarity refers to the resemblance of multiple tasks in their inputs, objectives, actions and rewards. Data-placement and data-migration tasks have different inputs and objectives (See §3.3.1). Prior works (yang2020multi, ; yu2020meta, ; teh2017distral, ; sun2020adashare, ; ammar2014automated, ; carroll2005task, ) show that a single RL agent cannot effectively learn optimal policies for multiple tasks with low task similarity because transferring the learning from one task to a dissimilar task can inhibit performance. Second, a single RL agent cannot perform two tasks concurrently. Since it performs data placement in the critical path of I/O request handling, a single agent cannot concurrently identify migration candidates, which impacts the timeliness of data migration. In §3.2, we quantitatively demonstrate that a single RL agent (SAPM) shows limited performance when handling both data placement and data migration.
We conclude that a single RL agent cannot effectively optimize both data placement and data migration in HSS, which motivates the need for a multi-agent RL approach.
@@ 3.4. Our Goal
Based on our observations and analyses in §3.2 and §3.3, we draw four conclusions. First, a holistic data-management technique is required to maximally utilize the potential of HSS. Second, RL-based techniques significantly improve performance over heuristic- and supervised-learning-based approaches. Third, given the low effectiveness of the naive combinations of existing techniques, the holistic data-management technique needs to coordinate data placement and data migration efficiently. Fourth, despite the high potential of RL for HSS data management, a single-agent-based approach has limited capability to optimize data placement and data migration simultaneously.
Our goal is to develop a new, holistic data-management technique for HSS that optimizes both data-placement and data-migration policies in a coordinated way to improve the overall HSS performance.
@@ 4. Harmonia
We design Harmonia as an online multi-agent RL technique for data management in HSS, which employs two autonomous RL agents: a data-placement agent and a data-migration agent. We implement Harmonia in the HSS management layer in the host operating system. In this section, we first describe the challenges of designing a multi-agent RL-based technique for HSS data management, and then explain Harmonia’s design and implementation in detail.
4.1. Challenges of Designing a Multi-Agent Technique for HSS
Designing a multi-agent RL-based data-management technique for HSS poses three key challenges. First, we need to design two autonomous RL agents to optimize the two different tasks of data placement and data migration (see §3.3.1). Second, we need to carefully formulate the reward structures of the two RL agents to enable coordination between the two agents (see §4.3) and improve HSS performance. Third, a multi-agent RL technique can increase computational complexity and storage overhead compared to a single-agent RL technique (see §4.6). We design Harmonia to address these challenges effectively.
While multi-agent RL itself is not a fundamentally new approach to system optimization, these prior techniques (e.g., (rl_opti_TACO_2017, ; qiu2022reinforcement, ; qiu2022simppo, ; qiu2020firm, ; mao2023multi, ; mlsys2023qiu, ; lu2022rlrp, ; jain2017cooperative, ; jain2017coordinated, ; shen2022multiagent, )) cannot be adapted directly for data management in HSS because they use (1) identical RL agents that focus on a single objective, and (2) joint-action space and shared reward functions for all RL agents.
@@ 4.2. Harmonia: RL Formulation
Fig. 5 shows the overview of Harmonia’s multi-agent RL framework. The key goal of the data-placement agent is to place incoming I/O requests in the best-fit storage device in the HSS, and that of the data-migration agent is to identify migration candidates from previously-placed pages in the HSS and determine their target storage device. We design the state, action, and reward of the two agents to enable coordination between them and maximize the overall HSS performance.
@@
State. Table 2 summarizes the state features for Harmonia’s two RL agents. We design a common set of seven state features because both agents consider the current HSS conditions and data characteristics (e.g., access frequency, migration interval) when making their decisions. Request type ($req\_type$) indicates a read or a write operation. Request size ($req\_size$) is the number of 4KiB pages in the current I/O request. Access interval ($acc\_intr$) is the number of page accesses between two accesses to the same page, while access frequency ($acc\_freq$) is the total number of accesses to the page. Fast-device capacity ($fast\_cap$) is the remaining free space in the fast storage device. Current device ($curr\_dev$) indicates the storage device where the page currently resides. Migration interval ($migr\_intr$) is the number of migrations performed between two migrations of the current page. The input features are represented as a 7-dimensional tuple (observation vector) for both agents:
@@
The size of the observation vector directly impacts Harmonia’s storage and computational overheads (see §4.6), so we optimize the observation vector size in two ways. First, we carefully select a limited set of highly relevant features. Second, we group the possible values of each input feature into a small number of bins (Table 2). For features with continuous values, we choose the number of bins based on characterization of evaluated workloads. For example, the request size can fall into any of eight size classes (e.g., $\leq$ 4 KiB, 8 KiB, 16 KiB, and so on) requiring three bits to represent this feature. In total, we use 32 bits for encoding the state representation.
@@
Action. The action space of Harmonia’s data-placement and data-migration agents depends on the number of storage devices in the HSS. The data-placement agent receives the state features for each incoming write request, and its action is to determine the best-fit storage device to place the I/O request data. The data-migration agent continuously looks at previously-placed pages in the HSS, and uses their state features to determine their target device for migration. Therefore, the two agents do not perform their actions in synchronization. In an HSS with two devices, (1) the data-placement agent can place data in either the fast or the slow device, and (2) the data-migration agent can migrate data to the fast or the slow device. The action sets of both agents are (1) small, which results in less computational complexity, and (2) easily extensible to multiple storage devices. We use a 4-bit encoding for the action to allow extensibility to more storage devices.
Reward. The reward structure is key to finding an optimal policy in RL. We design the reward structures of Harmonia’s data-placement and data-migration agents based on two key factors. First, since we focus on improving HSS performance, we use I/O request latencies in our reward structure for both agents. I/O request latencies capture the internal state (e.g., queueing delays, buffer dependencies, bandwidth utilization, effects of garbage collection, read/write latencies, non-blocking operations, error handling latencies) of storage devices in the HSS. Second, we design the reward structures to enable coordination between the two agents.
For every data-placement action at time step $t$, the data-placement agent receives a reward $R_{placement}$ from the environment at time step $t+1$. The reward for the placement agent is inversely proportional to the I/O request latency:
@@
We design a different reward structure for the migration agent since its objective is different from that of data placement (see §3.3.1). The data-migration agent’s objective is to maximize the long-term performance of the HSS. Unlike the immediate impact of data placement on HSS performance, the impact of data migrations may be delayed. This is due to two key reasons: (1) the application may not access the migrated data to the fast storage device immediately, and (2) the application may not immediately leverage the free space created in the fast storage device by cold data migration. Hence, we provide a delayed reward to the migration agent to capture the long-term impact of its decisions. The reward for the migration agent is:
@@
where $t$, $n$, $L_{i}$, $P_{migr_{t}}$ and $x$ represent the current time step, the number of incoming I/O requests whose data-placement latencies are considered, the latency of I/O request $i$, a penalty based on migration and access intervals, and the number of pages migrated before providing a reward, respectively. This reward is based on the average data-placement latencies (i.e., $\frac{\sum L_{i}}{n}$) for $n$ I/O requests that arrive after $x$ migration candidates are moved to their target devices, and serves as feedback for the migration decisions made previously. To avoid repeated migrations of the same page (ping-pong migrations), we add a small penalty $P_{migr_{t}}$ whose value is inversely proportional to the average migration and access intervals (See Table 2) of $x$ pages. Based on empirical analysis (see §6.1 for more details), we set the value of $n$ to 50 incoming requests and $x$ to 10, the size of the migration queue. We use a half-precision floating-point (16-bit) representation for the reward structure of both agents to reduce the storage overhead of training data. Our design of the reward structures of the two agents addresses the second key challenge described in §4.1.

### v2

@@ 3.3. Prior Approaches to HSS Data Management
Data-Placement Techniques. We evaluate two state-of-the-art HSS data-placement techniques: CDE (Cold Data Eviction) (matsui2017design, ) and Sibyl (singh2022sibyl, ). CDE is a heuristic data-placement technique that places hot or random (cold or sequential) data in the fast (slow) device, by determining the hotness and randomness of incoming data based on past access frequency and request size, respectively. Sibyl (singh2022sibyl, ) uses an RL agent to place data in the best-fit HSS storage device based on multiple features of I/O requests (e.g., request type and size) and HSS conditions (e.g., device capacity, I/O latency). When the fast device becomes full, both techniques evict the least-recently-used (LRU) data before placing new I/O request data in the fast device.
@@
Data-Migration Techniques. We evaluate two state-of-the-art HSS data-migration techniques, K-SVM (shetti2019machine, ) and RNN-HSS (doudali2019kleio, ), which have three aspects in common. First, they both leverage supervised learning to identify the target device for migration of each stored page within the HSS based on the page’s access frequency. Second, they periodically determine migration candidates (i.e., pages) and perform data migration at fixed intervals (e.g., migration is performed after every 1000 incoming requests) on the critical path of I/O request handling. Third, they perform data placement for incoming I/O requests using a fixed heuristic policy (e.g., K-SVM places incoming requests in the fast device until its capacity reaches 100% (shetti2019machine, )). The key difference between the two techniques is in the supervised-learning method used; K-SVM uses a k-means-assisted support vector machine (SVM) to classify stored pages as belonging to a storage device based on their past access frequencies; RNN-HSS uses recurrent neural networks (RNN) to predict a page’s future access frequency, and migrate the page to its target device.
@@
Extended Data-Management Techniques. Since there is no existing technique that optimizes both data placement and data migration together, we evaluate four extended techniques to explore the design space for a holistic data-management technique. For the first two techniques, Sibyl+K-SVM and Sibyl+RNN-HSS, we simply combine Sibyl, the current best-performing data-placement technique, with the prior data-migration techniques explained above.
@@
We design the third extended data-management technique, CDE+RL-Migr, [^2] The RL-based data-migration technique in CDE+RL-Migr is a contribution of this work. We design this as there are no prior RL-based HSS data-migration techniques. by combining CDE, a heuristic data-placement technique, with an RL-based data-migration approach. The RL-based data-migration approach uses the same design as the data-migration agent proposed in Harmonia, including the state features and the reward structure (See §4.2).
@@
In the fourth extended technique, SAPM (Single Agent for Placement and Migration), we design a single RL agent to perform both data placement and data migration to analyze the effectiveness of a single RL agent at HSS data management. SAPM performs a placement action for every incoming I/O request, and focuses on data migration only during system idle times to avoid interference with the placement task (which is on the critical path of I/O request handling). We provide the same state features used in Harmonia (See §4.2) to SAPM and tune the hyperparameters to achieve the best performance possible with a single RL agent. We provide SAPM’s RL agent (1) an immediate reward for every placement action (based on the reward of Harmonia’s data-placement agent in Equation 2), and (2) a deferred reward for the migration actions (based on the reward structure of Harmonia’s data-migration agent in Equation 3). We provide background in §2.3 on why a single RL agent cannot effectively optimize two different tasks.
@@ 3.4. Effectiveness of Prior Techniques
Methodology. We evaluate eight prior data-management techniques on a real system with two HSS configurations, performance- and cost-optimized, on 17 data-intensive workloads from 3 benchmark suites (see Table 4). See §5 for our evaluation methodology.
Performance Results. Fig. 3 shows the performance of four prior and four extended HSS data-management techniques on (a) performance-optimized and (b) cost-optimized HSS. We compare their average request latencies against two ideal approaches, Fast-Only and Oracle. First, in Fast-Only, we assume that the fast storage device is large enough to accommodate the entire workload dataset. Second, Oracle (meswani2015heterogeneous, ) makes optimal data-placement and migration decisions based on complete knowledge of future I/O access patterns of the entire workload. We assume that Oracle performs data migration during system idle times, which does not add any latency overheads. All values in Fig. 3 are normalized to Fast-Only.
@@
We make five key observations from Fig. 3. First, all prior techniques exhibit significant performance gaps from Oracle for every tested HSS and workload. The best-performing technique, Sibyl, only achieves 49.9% (41.4%) of Oracle’s performance in performance- (cost-) optimized HSS, which strongly implies the need for holistic data management in HSS. Sibyl’s RL-based data-placement technique can adapt to workload and system changes, but its heuristic data-eviction policy on the critical path lowers its performance.
@@
Second, Sibyl outperforms all other prior techniques, improving the average performance over CDE, K-SVM, and RNN-HSS by 20.7% (14.5%), 24.7% (31.6%), and 19.3% (19.2%), respectively, in the performance- (cost-) optimized HSS. Sibyl outperforms CDE, K-SVM and RNN-HSS because they use heuristic- or supervised-learning-based approaches, which cannot easily adapt to changes in workload access patterns and HSS conditions.
@@
Third, Sibyl+K-SVM and Sibyl+RNN-HSS underperform compared to Sibyl. A naive combination of prior approaches performs poorly as the placement and migration policies are oblivious to each other’s decisions. This leads to (1) conflicting decisions from the two policies, and (2) delayed data-placement policy convergence.
@@
Fourth, CDE+RL-Migr shows lower performance than Sibyl by 4.6% (15.6%) on average in performance- (cost-) optimized HSS. CDE+RL-Migr shows comparable performance to Sibyl in read-intensive workloads (i.e., RocksDB, MLPerf) because the RL-based migration agent proactively prefetches frequently-read data to the fast device, which Sibyl does not do. However, it underperforms in write-intensive workloads (i.e., SYSTOR17) because CDE’s heuristic data-placement policy places most incoming data in the fast storage device without considering device characteristics, resulting in frequent migrations. CDE+RL-Migr’s performance suggests that an RL-based data-migration policy can benefit performance if combined with an adaptive placement policy.
@@
Fifth, SAPM shows lower performance compared to Sibyl by 18.8% (27.7%) on average in the performance- (cost-) optimized HSS, even though SAPM uses RL for both data placement and migration. SAPM’s limited performance is due to two key reasons. (1) SAPM’s single RL agent is unable to learn an optimal policy for data placement and data migration concurrently, and (2) SAPM cannot identify migration candidates in a timely manner because it focuses on the data-migration task only when there are no incoming I/O requests. We provide a detailed analysis on why a single RL agent cannot optimize two different tasks in §3.2.
@@ 3.5. Our Goal
Based on our observations and analyses in §3.1, §3.2, §3.3, and §3.4, we conclude that a multi-agent RL approach’s ability to optimize different tasks concurrently alleviates the limitations of a single-agent RL approach.
Our goal is to design a holistic HSS data-management technique that optimizes both data-placement and data-migration policies using a multi-agent RL approach to improve the overall HSS performance.
@@ 4. Harmonia
We design Harmonia as an online multi-agent RL technique for HSS data management with two autonomous RL agents: a data-placement agent and a data-migration agent. We implement Harmonia in the HSS management layer in the host operating system. In this section, we first describe the challenges of designing a multi-agent RL-based technique for HSS data management, and then explain Harmonia’s design and implementation in detail.
4.1. Challenges of Designing a Multi-Agent HSS Technique
Designing a multi-agent RL-based HSS data-management technique poses three key challenges. First, we need to design two autonomous RL agents to optimize the different tasks of data placement and migration (see §3.2). Second, we need to formulate the reward structures of the two RL agents to enable coordination between them (see §4.3) and improve HSS performance. Third, having multiple RL agents can increase computational complexity and storage overhead (see §4.6). While multi-agent RL itself is not a fundamentally new approach to system optimization, prior techniques (e.g., (rl_opti_TACO_2017, ; qiu2022reinforcement, ; qiu2022simppo, ; qiu2020firm, ; mao2023multi, ; mlsys2023qiu, ; lu2022rlrp, ; jain2017cooperative, ; jain2017coordinated, ; shen2022multiagent, )) cannot be adapted directly for HSS data management because they use (1) identical RL agents with a single objective, and (2) joint-action space and shared reward functions for all RL agents.
@@ 4.2. Harmonia: RL Formulation
Fig. 4 shows the overview of Harmonia’s multi-agent RL framework. The key goal of the data-placement agent is to place incoming I/O requests in the best-fit storage device, and that of the data-migration agent is to identify migration candidates from previously-placed pages and determine their target storage device. We design the state, action, and reward of the two agents to enable coordination between them and maximize the overall HSS performance.
@@
State. Table 1 summarizes the state features for Harmonia’s two RL agents. We design a common set of seven state features because both agents consider the current HSS conditions and data characteristics when making their decisions. The state features are represented as a 7-dimensional tuple (observation vector) for both agents:
@@
The observation vector size directly impacts Harmonia’s storage and computational overheads (see §4.6), so we (1) carefully select a limited set of highly relevant features, and (2) group state feature values into a small number of bins (Table 1) based on workload characterization. For example, request size can fall into eight size classes (e.g., $\leq$ 4 KiB, 8 KiB, 16 KiB) requiring three bits to represent. In total, we use 32 bits for encoding the state representation.
@@
Action. The action space of Harmonia’s agents depends on the number of devices in the HSS. The data-placement agent receives the state features of incoming write requests, and its action is to determine the best-fit device to place request data. The data-migration agent continuously looks at the state features of previously-placed pages to determine their target device for migration. Therefore, the two agents do not perform their actions in synchronization. In an HSS with two devices, (1) the data-placement agent can place data in the fast or the slow device, and (2) the data-migration agent can migrate data to the fast or the slow device. The action sets of both agents are small and extensible, resulting in less computational complexity and easier multi-device scalability. We encode actions with 4 bits, allowing extensibility to more storage devices.
Reward. The reward structure is key to finding an optimal policy in RL. We design the reward structures of Harmonia’s agents based on two key factors. First, since we focus on HSS performance, we use I/O request latencies in our reward structure for both agents. I/O request latencies capture the internal state (e.g., queueing delays, buffer dependencies, bandwidth utilization, effects of garbage collection, read/write latencies, non-blocking operations, error handling latencies) of storage devices in the HSS. Second, we design the reward structures to enable coordination between the two agents.
For every placement action at time step $t$, the data-placement agent receives a reward $R_{placement}$ from the environment at time step $t+1$ that is inversely proportional to the I/O request latency:
@@
The data-migration agent’s objective is to maximize long-term HSS performance, which is different from the placement agent’s objective (see §3.2). Unlike the immediate impact of data placement on HSS performance, the impact of data migrations (i.e., prefetching frequently-accessed data to the fast device and evicting cold data to the slow device) may be delayed due to two key reasons: (1) the application may not access the prefetched data immediately, and (2) the application may not immediately leverage the free space created in the fast device by cold data migration. Hence, we provide a delayed reward to the migration agent to capture the long-term impact of its decisions. The reward for the migration agent is:
@@
where $t$ is the current time step, $n$ is the number of incoming I/O requests whose data-placement latencies are considered, $L_{i}$ is the latency of I/O request $i$, $P_{migr_{t}}$ is a penalty based on migration and access intervals, and $x$ is the number of pages migrated before providing a reward. This reward is based on the average data-placement latencies (i.e., $\frac{\sum L_{i}}{n}$) for $n$ I/O requests that arrive after $x$ migration candidates are moved to their target devices, and serves as feedback for the migration decisions made previously. To avoid repeated migrations of the same page (ping-pong migrations), we add a small penalty $P_{migr_{t}}$ whose value is inversely proportional to the average migration and access intervals (See Table 1) of $x$ pages. Based on empirical analysis (see §6.1), we set the value of $n$ to 50 and $x$ to 10, the size of the migration queue. We use a half-precision floating-point (16-bit) representation for the reward structure of both agents to reduce the storage overhead of training data.